If 50 ml of 0.3 F HAc is mixed with 100 ml of 0.1 F KOH, what is the hydronium-ion concentration of the resulting solution?

What is the pH of the solution?


1
Expert's answer
2022-03-09T11:19:04-0500


50 mL = 0.05 L

0.3 * 0.05 = 0.015 (HAc)

100 mL = 0.1 L

0.1 * 0.1 = 0.01 (KOH)

Δ = 0.015 – 0.01 = 0,005 (HAc)

CH3COOH + HOH = H3O+ + CH3COO-

Ka = 1.8*10-5

 = 1.8*10-5

x = 0.00029

pH = - log10(0.00029) = 3.5